Solve using the substitution method;
2x + y = 5
x = y + 7

Answer by TimothyLamb(4379)

2x + y = 5
x = y + 7
---
put the system of linear equations into standard form
---
2x + y = 5
x - y = 7
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 4
y = -3
---
now by substitution:
---
2x + y = 5
x = y + 7
---
2(y + 7) + y = 5
2y + 14 + y = 5
3y = -9
y = -3
x = -3 + 7
x = 4
---
solution by substitution:
x = 4
y = -3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---